Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2025 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 40     Entry time: 03 Apr 2025, 07:34
Author: Harald Sinn 
Group: PRC 
Subject: SA1 status 

7:20 FXE reports a problem with laser synchronisation (yellow warning light and could not change delay). BKR informs the DESY laser sync group. 

7:51 DESY sync group reported that there was a CPU failing about 6 am this morning which led to the yellow warning light. However, synchronisation was very likely all the time ok withing 5-10fs. The CPU was now reset and everything should be wokring now. 

ELOG V3.1.4-7c3fd00